# Algebraic Independence of Special Points on Shimura Varieties

### v1

@@ 4.1. Proof of the main theorem over $\overline{\mathbb{Q}}$
Let the image of $R$ under these transformations be $R^{\prime}\subset F_{2}\times F_{T_{0}}$. By applying the point counting theorem if necessary, we may take $R^{\prime}$ to be connected and semialgebraic. Since the projection of $R$ to $\mathbb{R}^{\ell}$ was non-constant, the projection of $R^{\prime}$ to $F_{T_{0}}$ is also non-constant. Moreover, we may assume that $T/T_{0}$ is a semi-abelian variety by expanding $T_{0}$ if necessary because the projection to the affine space $\mathbb{G}_{a}$ is constant ($b_{l}=0$). By applying the Ax–Schanuel theorem (Theorem 2.5) to the Zariski closure of $(F_{1}\times R^{\prime})\times(q\times\exp)(F_{1}\times R^{\prime})$, we get that $(q\times\exp)(F_{1}\times R^{\prime})\subset S\times T/T_{0}$ is contained in a proper weakly special subvariety. Since $R^{\prime}$ contains preimages of special points, its image must lie in a proper special subvariety $S^{\prime}\times T^{\prime}\subset S\times T/T_{0}$.
@@
Let $V^{\prime}\subset S\times T/T_{0}$ be the image of $V$ under the quotient map $T\to T/T_{0}$. By construction, we have that $(q\times\exp)(F_{1}\times R^{\prime})\subset V^{\prime}$, and so $V^{\prime}$ contains a proper special subvariety $S^{\prime}\times T^{\prime}$. However, the map $V\to S$ is finite, and hence $V^{\prime}\to S$ is also finite. Hence, $T^{\prime}$ is a single point, meaning that $R$ was originally of the form $R^{\prime\prime}\times\{g\}$, with $g$ corresponding to a torsion translation of $T_{0}$, which by abuse of notation we will also denote $T_{0}$. Let $\tilde{T_{0}}\subset F_{T}$ be the preimage of $T_{0}$. Applying Ax–Schanuel to the Zariski closure of
$U\coloneqq(F_{1}\times R^{\prime\prime}\times\tilde{T_{0}})\times(q(F_{1}\times R^{\prime\prime})\times T_{0})\cap V,$
@@
we get that $q(F_{1}\times R^{\prime\prime})\times T_{0}$ is contained in a proper special subvariety of $V$, which must be of the form $S^{\prime}\times T_{0}\subset V$. This shows that our original exemplary component $W$ was not exemplary because there is a larger Shimura variety $S^{\prime}$ properly containing $\pi_{S}(W)$ whose $V$ image lies within $T_{0}=T^{\prime}$. ∎
@@ 4.2. From $\overline{\mathbb{Q}}$ to $\mathbb{C}$
Let $F$ be the finitely generated subfield over $\mathbb{Q}$ which $S,T,\operatorname{End}(T)$, and $V$ are all defined. Then, there exists a quasi-projective geometrically irreducible variety $Z$ over a number field $K$ such that $F$ is the function field of $Z$, and $S$ is defined over $K$. By taking a Zariski open subset of $Z$, we may assume that $T$ extends to an abelian scheme $\mathcal{T}$ over $Z$ and $V$ extends to a variety $\mathcal{V}$ that is flat over $Z$. Choose a generic closed point $z_{0}\in Z(\mathbb{C})$ such that the field gotten by localizing at $z_{0}$ is isomorphic to $F$. Choose a simply connected Euclidean open neighborhood $U\ni z_{0}$ of $z_{0}$ so that the homology of $\mathcal{T}$ can be trivialized over $U$, and so we have $\mathcal{T}_{U}\cong T\times U$ as analytic varieties.
@@
Suppose for the sake of contradiction that there are infinitely many exemplary subvarieties of $V\cong\mathcal{V}_{z_{0}}$. By Lemma 4.2, special subvarieties that arise as projections of the exemplary subvarieties of $V$ to $S$ come from finitely many splittings of special subvarieties of $S$. So, there is a splitting of sub-Shimura datum $(H_{1},X_{1})\times(H_{2},X_{2})$ with image $S_{1}\times S_{2}\subset S$ and infinitely many special points $p_{i}\in S_{2}$ such that there are exemplary subvarieties $W_{i}$ of $V$ mapping surjectively to $S_{1}\times\{p_{i}\}$. By the André–Oort conjecture, proven by [PST+21], the Zariski closure of these $S_{1}\times\{p_{i}\}$ is a finite union $R_{1},\dots,R_{n}\subset S$ of special subvarieties of $S$. For each $i$, let $T_{i}\subset T$ be the smallest special subvariety containing $\pi_{T}(W_{i})\subset\mathcal{T}_{z_{0}}$, and we extend $T_{i}$ to a family $\mathcal{T}_{i}$ over $U$. For every $z\in U$, we have that the $\mathcal{V}_{z}$ image of $S_{1}\times\{p_{i}\}$ lies within $T_{i,z}$. However, for each $z\in U(\overline{\mathbb{Q}})$, Theorem 4.4 says that there are only finitely many exemplary components and hence the image of each $R_{i}$ must also lie within a proper special subvariety of $T$.
@@
The $\mathcal{V}$-images of each $R_{i}$ gives a family of subvarieties $\mathcal{R}$ of $\mathcal{T}$ such that for each $z\in U(\overline{\mathbb{Q}})$, the fiber $\mathcal{R}_{z}$ lies within a proper special subvariety of $\mathcal{T}_{z}$. We claim that this holds at $z_{0}$ as well. By replacing $\mathcal{R}$ with a $g$ self-sum, where $g=\dim T$, we may assume the $\mathcal{R}$ is a coset of an abelian subscheme of $\mathcal{T}$ and we still have that $\mathcal{R}_{z}$ lies within a proper special subvariety of $T$. By quotienting by the identity component of $\mathcal{R}$, we may assume that $\mathcal{R}$ is finite over $U$ and after finite base change that $\mathcal{R}$ is a section of $U$. Applying the Main Theorem of [Mas89] and the extension to semi-abelian varieties given in Section $5$ of loc. cit. to ×$\mathcal{R}\times\tau_{1}\mathcal{R},\dots,\tau_{k}\mathcal{R}\subset T^{k+1}$, we get that $\mathcal{R}_{z}$ must lie within a proper special subvariety of $\mathcal{T}_{z}$. ∎
@@ 5. Applications of the main theorem
We may assume that $S$ is a connected Shimura variety whose Shimura datum $(G,X)$ is such that $G$ is a product of simple groups and that $S^{\prime}$ is also a connected Shimura variety whose Shimura datum is $(G^{\prime},X^{\prime})$. By viewing $S^{\prime}$ as the orbit of a point under the action of a subgroup of $G^{n}$, it suffices to show that if $G^{\prime}\subset G_{1}\times\cdots G_{n}$ is a proper subgroup of a product of simple groups, then there exist $i,j$ such that $\pi_{ij}(G^{\prime})\subset G_{i}\times G_{j}$ is a proper subgroup.
@@
Let $S,T,V$ be as in Corollary 1.3. Fix $n\in\mathbb{N}$. There exists a proper Zariski closed subset $S^{\prime}\subset S^{2}$ such that for any $n$ points $s_{1},\dots,s_{n}\in S$ such that $(s_{i},s_{j})\not\in S^{\prime}$ for any pair $i\neq j$, then any $V$-images $t_{1},\dots,t_{n}$ for $s_{1},\dots,s_{n}$ are linearly independent.
@@
If $x_{1},\cdots,x_{n}$ are linearly dependent in $T$ then by Corollary 5.2, there must exist some $i\neq j$ such that $(s_{i},s_{j})$ lies within a finite set of proper special subvarieties of $S^{2}$. So we need to describe the special subvarieties in $\mathcal{A}_{g}\times\mathcal{A}_{g}$.
@@
By applying Corollary 5.2 to $V^{r+1}$ inside of $S^{r+1}\times T^{r+1}$, we obtain that there exists finitely many proper special subvarieties $S_{n}$ of $S^{2}$ such that for any $r+1$ points $s_{1},\dots,s_{r+1}\in S$ with $(s_{i},s_{j})\not\in\bigcup_{n}S_{n}$ for all pairs $i,j$, any $V$-images of the $s_{i}$ are linearly independent. Since the Shimura datum is simple, the argument in the proof of Corollary 1.4 gives that either the projection of $S_{n}$ on one factor is a proper special subvariety or $S_{n}$ is a correspondence on $S^{2}$. Let $S^{\prime}$ be the union of all of the proper projections and let $D$ be the maximum degree of the projections of $S_{n}$ to a factor of $S$ when $S_{n}$ is a correspondence. Then, there exists an $N=N(r+1,D)$ such that given $s_{1},\dots,s_{N}\in S\backslash S^{\prime}$ special, we can find a subset $s_{1}^{\prime},\dots,s_{r+1}^{\prime}$ such that no $(s_{i}^{\prime},s_{j}^{\prime})$ lies in a $S_{n}$, and thus their $V$-images are linearly independent. Thus, outside of $S^{\prime}\subset S$, there are less than $N$ other special points of $S$ whose $V$-images lie in $\Gamma^{\prime}$. Both $S^{\prime}$ and $N$ depend on $S,T,V,r$, and are independent of $\Gamma$, proving the theorem. ∎

### v2

@@ 4.1. Proof of the main theorem over $\overline{\mathbb{Q}}$
Let the image of $R$ under these transformations be $R^{\prime}\subset F_{2}\times F_{G_{0}}$. By applying the point counting theorem if necessary, we may take $R^{\prime}$ to be connected and semialgebraic. Since the projection of $R$ to $\mathbb{R}^{\ell}$ was non-constant, the projection of $R^{\prime}$ to $F_{G_{0}}$ is also non-constant. Moreover, we may assume that $G/G_{0}$ is a semi-abelian variety by expanding $G_{0}$ if necessary because the projection to the affine space $\mathbb{G}_{a}$ is constant ($b_{l}=0$). By applying the Ax–Schanuel theorem (Theorem 2.5) to the Zariski closure of $(F_{1}\times R^{\prime})\times(q\times\exp)(F_{1}\times R^{\prime})$, we get that $(q\times\exp)(F_{1}\times R^{\prime})\subset S\times G/G_{0}$ is contained in a proper weakly special subvariety. Since $R^{\prime}$ contains preimages of special points, its image must lie in a proper special subvariety $S^{\prime}\times G^{\prime}\subset S\times G/G_{0}$.
@@
Let $V^{\prime}\subset S\times G/G_{0}$ be the image of $V$ under the quotient map $G\to G/G_{0}$. By construction, we have that $(q\times\exp)(F_{1}\times R^{\prime})\subset V^{\prime}$, and so $V^{\prime}$ contains a proper special subvariety $S^{\prime}\times G^{\prime}$. However, the map $V\to S$ is finite, and hence $V^{\prime}\to S$ is also finite. Hence, $G^{\prime}$ is a single point, meaning that $R$ was originally of the form $R^{\prime\prime}\times\{g\}$, with $g$ corresponding to a torsion translation of $G_{0}$, which by abuse of notation we will also denote $G_{0}$. Let $\tilde{G_{0}}\subset F_{G}$ be the preimage of $G_{0}$. Applying Ax–Schanuel to the Zariski closure of
$U\coloneqq(F_{1}\times R^{\prime\prime}\times\tilde{G_{0}})\times(q(F_{1}\times R^{\prime\prime})\times G_{0})\cap V,$
@@
we get that $q(F_{1}\times R^{\prime\prime})\times G_{0}$ is contained in a proper special subvariety of $V$, which must be of the form $S^{\prime}\times G_{0}\subset V$. This shows that our original exemplary component $W$ was not exemplary because there is a larger Shimura variety $S^{\prime}$ properly containing $\pi_{S}(W)$ whose $V$ image lies within $G_{0}=G^{\prime}$. ∎
@@ 4.2. From $\overline{\mathbb{Q}}$ to $\mathbb{C}$
Let $F$ be the finitely generated subfield over $\mathbb{Q}$ which $S,G,\operatorname{End}(G)$, and $V$ are all defined. Then, there exists a quasi-projective geometrically irreducible variety $Z$ over a number field $K$ such that $F$ is the function field of $Z$, and $S$ is defined over $K$. By taking a Zariski open subset of $Z$, we may assume that $G$ extends to an abelian scheme $\mathcal{G}$ over $Z$ and $V$ extends to a variety $\mathcal{V}$ that is flat over $Z$. Choose a generic closed point $z_{0}\in Z(\mathbb{C})$ such that the field gotten by localizing at $z_{0}$ is isomorphic to $F$. Choose a simply connected Euclidean open neighborhood $U\ni z_{0}$ of $z_{0}$ so that the homology of $\mathcal{G}$ can be trivialized over $U$, and so we have $\mathcal{G}_{U}\cong G\times U$ as analytic varieties.
@@
Suppose for the sake of contradiction that there are infinitely many exemplary subvarieties of $V\cong\mathcal{V}_{z_{0}}$. By Lemma 4.2, special subvarieties that arise as projections of the exemplary subvarieties of $V$ to $S$ come from finitely many splittings of special subvarieties of $S$. So, there is a splitting of sub-Shimura datum $(H_{1},X_{1})\times(H_{2},X_{2})$ with image $S_{1}\times S_{2}\subset S$ and infinitely many special points $p_{i}\in S_{2}$ such that there are exemplary subvarieties $W_{i}$ of $V$ mapping surjectively to $S_{1}\times\{p_{i}\}$. By the André–Oort conjecture, proven by [PST+21], the Zariski closure of these $S_{1}\times\{p_{i}\}$ is a finite union $R_{1},\dots,R_{n}\subset S$ of special subvarieties of $S$. For each $i$, let $G_{i}\subset G$ be the smallest special subvariety containing $\pi_{G}(W_{i})\subset\mathcal{G}_{z_{0}}$, and we extend $G_{i}$ to a family $\mathcal{G}_{i}$ over $U$. For every $z\in U$, we have that the $\mathcal{V}_{z}$ image of $S_{1}\times\{p_{i}\}$ lies within $G_{i,z}$. However, for each $z\in U(\overline{\mathbb{Q}})$, Theorem 4.4 says that there are only finitely many exemplary components and hence the image of each $R_{i}$ must also lie within a proper special subvariety of $G$.
@@
The $\mathcal{V}$-images of each $R_{i}$ gives a family of subvarieties $\mathcal{R}$ of $\mathcal{G}$ such that for each $z\in U(\overline{\mathbb{Q}})$, the fiber $\mathcal{R}_{z}$ lies within a proper special subvariety of $\mathcal{G}_{z}$. We claim that this holds at $z_{0}$ as well. By replacing $\mathcal{R}$ with a $g$ self-sum, where $g=\dim G$, we may assume the $\mathcal{R}$ is a coset of an abelian subscheme of $\mathcal{G}$ and we still have that $\mathcal{R}_{z}$ lies within a proper special subvariety of $G$. By quotienting by the identity component of $\mathcal{R}$, we may assume that $\mathcal{R}$ is finite over $U$ and after finite base change that $\mathcal{R}$ is a section of $U$. Applying the Main Theorem of [Mas89] and the extension to semi-abelian varieties given in Section $5$ of loc. cit. to ×$\mathcal{R}\times\tau_{1}\mathcal{R},\dots,\tau_{k}\mathcal{R}\subset G^{k+1}$, we get that $\mathcal{R}_{z}$ must lie within a proper special subvariety of $\mathcal{G}_{z}$. ∎
@@ 5. Applications of the main theorem
We may assume that $S$ is a connected Shimura variety whose Shimura datum $(G_{S},X)$ is such that $G_{S}$ is a product of simple groups and that $S^{\prime}$ is also a connected Shimura variety whose Shimura datum is $(G^{\prime},X^{\prime})$. By viewing $S^{\prime}$ as the orbit of a point under the action of a subgroup of $G_{S}^{n}$, it suffices to show that if $G^{\prime}\subset G_{1}\times\cdots G_{n}$ is a proper subgroup of a product of simple groups, then there exist $i,j$ such that $\pi_{ij}(G^{\prime})\subset G_{i}\times G_{j}$ is a proper subgroup.
@@
Let $S,G,V$ be as in Corollary 1.3. Fix $n\in\mathbb{N}$. There exists a proper Zariski closed subset $S^{\prime}\subset S^{2}$ such that for any $n$ points $s_{1},\dots,s_{n}\in S$ such that $(s_{i},s_{j})\not\in S^{\prime}$ for any pair $i\neq j$, then any $V$-images $g_{1},\dots,g_{n}$ for $s_{1},\dots,s_{n}$ are linearly independent.
@@
If $x_{1},\cdots,x_{n}$ are linearly dependent in $G$ then by Corollary 5.2, there must exist some $i\neq j$ such that $(s_{i},s_{j})$ lies within a finite set of proper special subvarieties of $S^{2}$. So we need to describe the special subvarieties in $\mathcal{A}_{g}\times\mathcal{A}_{g}$.
@@
By applying Corollary 5.2 to $V^{r+1}$ inside of $S^{r+1}\times G^{r+1}$, we obtain that there exists finitely many proper special subvarieties $S_{n}$ of $S^{2}$ such that for any $r+1$ points $s_{1},\dots,s_{r+1}\in S$ with $(s_{i},s_{j})\not\in\bigcup_{n}S_{n}$ for all pairs $i,j$, any $V$-images of the $s_{i}$ are linearly independent. Since the Shimura datum is simple, the argument in the proof of Corollary 1.4 gives that either the projection of $S_{n}$ on one factor is a proper special subvariety or $S_{n}$ is a correspondence on $S^{2}$. Let $S^{\prime}$ be the union of all of the proper projections and let $D$ be the maximum degree of the projections of $S_{n}$ to a factor of $S$ when $S_{n}$ is a correspondence. Then, there exists an $N=N(r+1,D)$ such that given $s_{1},\dots,s_{N}\in S\backslash S^{\prime}$ special, we can find a subset $s_{1}^{\prime},\dots,s_{r+1}^{\prime}$ such that no $(s_{i}^{\prime},s_{j}^{\prime})$ lies in a $S_{n}$, and thus their $V$-images are linearly independent. Thus, outside of $S^{\prime}\subset S$, there are less than $N$ other special points of $S$ whose $V$-images lie in $\Gamma^{\prime}$. Both $S^{\prime}$ and $N$ depend on $S,G,V,r$, and are independent of $\Gamma$, proving the theorem. ∎